A5% solution of can sugar is isotonic with 1% of a solution of an unknown solute the molar mass of unknown solute in g/mol is

Grade:12

1 Answers

ISHITAL PAL
5 years ago
since both the solutions are isotonic thus their osmotic pressure will be equal....hence C1RT=C2RT
Now R and T will be cancelled out,
C1=C2
where C1 is the concentration of cane sugar while C2 is the concentration of unknown solute,
No. of moles of cane sugar÷ volume(in L.) = no. Of moles of unknown solute ÷ volume(L ),
Volumes will be equal(100 L. Each)and thus cancelled out.
Now, 5/342 = 1/ x
x= 342÷5= 68.4
Hence the molecular mass of unknown solute will be 68.4😄
 
b